# Can LLMs Replace Manual Annotation of Software Engineering Artifacts?

### v1

@@ IV-A RQ1: Level of Agreement between Humans and Models
TABLE I: Summary of Inter-Rater Agreements. Three Best Performing Models (i.e., Claude, Gemini and GPT-4) have been Used to Report the Agreement in this Table.
@@
Figure 1: Inter-Rater Agreement (Krippendorff’s Alpha) for Code Summarization Accuracy and Similarity. For Adequacy and Conciseness, we have similar plots to Accuracy; thus, we omitted them from the paper due to page constraints.
@@
Figure 4: Inter-Rater Agreement for Semantic Similarity Evaluation Criteria ‘Goals’. For other criteria (‘Operations’ and ‘Effects’), we have very similar heatmaps; thus, they were omitted from paper.
@@
Figure 5: Inter-Rater Agreement for Static Analysis Warning.
We note that human-model agreement varies from task to task. While it resembles human-human agreement values for some tasks (code summarization, name-value inconsistencies, semantic similarity), it is lower for other tasks (causal relation detection and static analysis warnings).
@@ IV-B RQ2: LLMs’ Applicability for a Specific Task
We observed that for a majority of tasks, human-model inter-rater agreement is very similar to human-human agreement, while for some tasks, it is lower. What about model-model inter-rater agreement, which does not require any human effort? Figure 6 shows that the mean model-model agreement of the top 3 models is positively correlated with the mean human-model agreement, with a Spearman correlation of 0.65 and a $p-value<0.05$. This suggests that higher model-model agreement is helpful in indicating good human-model agreement; this observation is potentially valuable in deciding whether to replace humans with models. There is an outlier (high model-model agreement but low human-model agreement) in the bottom-right corner, but for this case (code summarization conciseness), human-human agreement is also low (0.24) suggesting that human ratings for this task per se are not consistent. Our results suggest that if multiple LLMs reach similar solutions independently, then LLMs are likely suitable for the annotation task.
Figure 6: Human-Model Inter-Rater Agreements is Positively Co-Related with Model-Model Inter-Rater Agreements with $p-value<0.05$. Human-Human agreement are shown in blue text.
@@ IV-C RQ3: LLMs’ Applicability for a Specific Sample
Figure 7: Inter-Rater Agreement with increasing % of Samples from GPT-4 for Code Summarization Accuracy and Similarity. The dotted line indicates the inter-human agreement for that task.
@@
Figure 10: Inter-Rater Agreement with increasing % of Samples from GPT-4 for Static Analysis Warning.
For causality and static analysis warning, we have low inter-model agreement (0.39 and 0.12). We have already discussed that we should expect low human-model agreement because they are correlated. Figures 9 & 10 show that the inter-rater agreement decreases as we increase the samples from the GPT-4 model, indicating that the model may not be applicable in this setup. However, using the confidence-based approach, we can still see that GPT-4 maintains similar inter-rater agreement, up to 60% for the causality dataset and 50% for the static analysis dataset, showing some promise for using GPT-4 for partial datasets. The confidence interval for these datasets up to that point completely overlaps with the confidence interval of the human-human rating. High-probability samples are likely to be more correct or align better with human preferences compared to other samples.
@@
TABLE II: % of Human-Effort Saved for One Rating and Overall Annotation Process.
@@ V Discussion & Limitations
Deciding Whether and When to Ask an LLM Our findings suggest that we can replace one human rating by a model for a significant number of samples (50%-100%). However, it appears risky to fully replace all humans, even though the probability of the model output at least somewhat indicates the quality of annotations. Based on our findings, we propose the process as given in Figure 11, which involves two steps: 1) Create few-shot examples (3–4 in our setup) to query multiple strong LLMs with all samples, and compute the model-model agreement. 2) If the agreement is high ($>0.5$ based on our data), one can safely replace one human rating per sample with an LLM-provided answer. Otherwise, selectively replace one rating only for those samples where the LLM gives a high-confidence answer.
Beyond saving human effort, our findings can also help extend a dataset by automatically labeling additional samples. We should emphasize that replacing more than one human can inflate inter-rater agreement because model-model agreements are much higher. This may not fulfill or reflect the original goals of the data annotations.
@@
Figure 13: Change of Fraction of Samples where GPT-4 Agress with Majority Voting for % of Samples Sorted Based on GPT-4 Model’s Output Probability.
@@
Limitations We consider only 10 tasks datasets and 6 models. It is quite difficult to find publicly available datasets where the ratings from individual annotators are still available; many artifacts contain only the final annotation after resolving disagreements. Even though we see pretty consistent results across the datasets and models, finding may vary with others. Only Krippendorff’s $\alpha$ was applicable to all studied datasets. Cohen’s $\kappa$ is another popular measure but is applicable to only two raters. Using few-shot learning means that some human effort is required initially. However, labeling 3–4 samples is much cheaper compared to labeling all the samples.
@@
VI Related Work
@@
ML and LLMs in Software Engineering Learning-based approaches and LLMs have been applied to various software development tasks [33], including code completion [34, 35, 36], test generation [37, 38, 39, 40], fuzzing [41, 42], automated repair [43, 44, 45, 46, 47] and coding agents [48, 49, 50, 51], and type prediction [52, 53, 54, 55]. We follow this trend, but deviate by not targeting a specific task, but by studying the potential for automating human-subject (manual) evaluations, using LLMs.
@@
LLMs to Complement or Replace Human Annotators Recent and concurrent work explores the idea of complementing or replacing human annotators with LLMs in domains beyond software engineering. Work in natural language processing (NLP) compares LLMs and human participants, typically recruited via crowd-working platforms, and report mixed results. While some studies show that adding LLM labels can improve the aggregated labels of a dataset [56] and sometimes outperform crowd-workers [57], others note that LLMs fail to accurately represent differences between demographic groups [58]. Others suggest to use LLMs to perform sub-tasks in a crowd-sourcing pipeline [59]. Bavaresco et al. [13] propose a benchmark of NLP datasets to compare human and LLM performance. We envision our work to serve as such a benchmark in software engineering. To the best of our knowledge, our paper is the first to carefully study LLMs for annotating software engineering artifacts, and propose actionable guidelines for when and how to use LLMs to create annotations.
@@
Collaborative Human-LLM Annotation Our suggested workflow (Section V) relates to work on human-LLM collaboration for data labeling [60]. Some papers [61, 12] propose to first query an LLM, and then let humans refine some of the LLM-provided labels. Instead, we find that inter-model agreement and LLM output probabilities provide an effective way of deciding which labeling tasks can be safely delegated to a model.
LLMs to Automate Research Tasks Our work fits into a broader stream of work on partially automating particular sub-tasks of performing academic research. Liang et al. [62] study to what extent LLMs can replicate empirical software engineering research by determining assumptions made in a study and by writing data processing code. While we share the idea of using LLMs to automate a research task, we focus on a different kind of task. Other work proposes to let LLMs generate open-ended questionnaire responses for HCI-style evaluations [63], and shows a set of LLMs can reproduce human-study results in economics, psycholinguistics, and social psychology [10]. We expect to see more work on delegating sub-tasks of conducting research to LLMs, e.g., for performing human-intensive evaluations and editing scientific texts.
@@
LLMs for Data Synthesis and Assessment Beyond using LLMs to partially automate evaluations that would otherwise be performed by humans only, others propose to use LLMs to synthesize additional data to train or fine-tune models for text annotation tasks [64, 65]. There is a recent related survey [66]. A key difference from our work is that training and fine-tuning datasets must be large-scale, but some amount of noise is acceptable, whereas annotations in human-subject studies are typically of smaller scale, but should have high confidence. LLMs have also been proposed as judges of outputs generated by other LLMs, either with a single LLM as the judge [67] or with a set of LLMs that discuss until reaching an agreement [68]. These efforts are primarily about determining human preference, i.e., tasks where different humans may legitimately disagree, whereas many annotation tasks in software engineering have an objectively correct answer that humans can eventually agree upon.
VII Conclusion
@@
In this paper, we investigate if LLM responses can substitute for human raters in software engineering annotation tasks, and how such substitutions may affect inter-rater agreement. We find that human-model agreements can be fairly consistent with human-agreements, for the majority of the settings we studied. We also find model-model agreement is a good indication of human-model agreement, suggesting that when powerful models trained on giant human corpora agree with each other, they tend to agree also with humans. Finally, we find that an LLM’s confidence (output probability) for a given sample output is a good indication of whether LLM output agrees with majority human rating, for that sample. We caution that the paper considers only discrete (multiple choice) LLM responses, not free-form annotations; furthermore, we have not studied issues of model output bias, nor issues of demographics. Finally, our scripts and datasets are made available anonymously at https://zenodo.org/doi/10.5281/zenodo.13146386

### v2

@@ IV-A RQ1: Level of Agreement between Humans and Models
TABLE I: Summary of inter-rater agreements. The three best performing models (Claude, Gemini, and GPT-4) have been used to report the agreement in this table.
@@
Figure 2: Inter-rater agreement (Krippendorff’s $\alpha$) for code summarization accuracy and similarity. Results for adequacy and conciseness are similar (omitted due to space).
@@
Figure 5: Inter-rater agreement for semantic similarity with evaluation criteria “Goals”. Other criteria (“Operations” and “Effects”) give similar heatmaps (omitted due to space).
@@
Figure 6: Inter-rater agreement for static analysis warnings.
We note that human-model agreement varies from task to task. While it resembles human-human agreement values for some tasks (code summarization, name-value inconsistencies, semantic similarity), it is lower for other tasks (causal relation detection and static analysis warnings). For some tasks, all reviewers labeled all the samples (e.g., static analysis warning and name-value inconsistency). For code summarization, the annotation load was equally distributed among the annotators. In code summarization, each pair of humans had 105 overlapping annotations. However, for causality, the load was not equally distributed. In some cases, there are no samples labeled by two individuals, and in those cases, the corresponding cell is empty.
@@ IV-B RQ2: LLMs’ Applicability for a Specific Task
We observed that for a majority of tasks, human-model inter-rater agreement is very similar to human-human agreement, while for some tasks, it is lower. What about model-model inter-rater agreement, which does not require any human effort? Figure 7 shows that the mean model-model agreement of the top 3 models is positively correlated with the mean human-model agreement, with a Spearman correlation of 0.65 ($p<0.05$). This suggests that higher model-model agreement is helpful in indicating good human-model agreement; this observation is potentially valuable in deciding whether to replace humans with models. There is an outlier (high model-model agreement but low human-model agreement) in the bottom-right corner, but for this case (code summarization conciseness), human-human agreement is also low (0.24) suggesting that human ratings for this task per se are not consistent. Our results suggest that if multiple LLMs reach similar solutions independently, then LLMs are likely suitable for the annotation task.
Figure 7: Human-model inter-rater agreements are positively correlated with model-model inter-rater agreements ($p<0.05$). Human-human agreements are shown in blue text.
@@ IV-C RQ3: LLMs’ Applicability for a Specific Sample
Figure 8: Inter-rater agreement with increasing % of samples from GPT-4 for code summarization accuracy and similarity. The dotted line indicates the inter-human agreement.
@@
Figure 11: Inter-rater agreement with increasing % of samples from GPT-4 for static analysis warnings.
For causality and static analysis warning, we have low inter-model agreement (0.39 and 0.12). We have already discussed that we should expect low human-model agreement because they are correlated. Figures 10 & 11 show that the inter-rater agreement decreases as we increase the samples from the GPT-4 model, indicating that the model may not be applicable in this setup. However, using the confidence-based approach, we can still see that GPT-4 maintains similar inter-rater agreement, up to 60% for the causality dataset and 50% for the static analysis dataset, showing some promise for using GPT-4 for partial datasets. The confidence interval for these datasets up to that point completely overlaps with the confidence interval of the human-human rating. High-probability samples are likely to be more correct or align better with human preferences compared to other samples.
@@
TABLE II: % of human effort saved for one rating and overall annotation process.
@@ V Discussion
Deciding Whether and When to Ask an LLM Our findings suggest that we can replace one human rating by a model for a significant number of samples (50%-100%). However, it appears risky to fully replace all humans, even though the probability of the model output at least somewhat indicates the quality of annotations. Based on our findings, we propose the process as given in Figure 12, which involves two steps: 1) Create few-shot examples (3–4 in our setup) to query multiple strong LLMs with all samples, and compute the model-model agreement. 2) If the agreement is high ($>0.5$ based on our data), one can safely replace one human rating per sample with an LLM-provided answer. Otherwise, selectively replace one rating only for those samples where the LLM gives a high-confidence answer. Beyond saving human effort, our findings can also help extend a dataset by automatically labeling additional samples. We should emphasize that replacing more than one human can inflate inter-rater agreement because model-model agreements are much higher. This may not fulfill or reflect the original goals of the data annotations.
@@
Figure 14: Change of fraction of samples where GPT-4 agrees with majority voting for % of samples sorted based on GPT-4 model’s output probability.
@@
VI Threats to Validity
We consider only 10 human-annotation tasks and 6 models. It is difficult to find publicly available datasets where the ratings from individual annotators are still available; many artifacts contain only the final annotation after resolving disagreements. Even though we see fairly consistent results across the datasets and models, finding may vary with others. Only categorical Krippendorff’s $\alpha$ was applicable to all studied datasets. For some of our tasks, ordinal Krippendorff’s $\alpha$ is also applicable (e.g., code summarization). However, we observe negligible change from using ordinal Krippendorff’s $\alpha$.
Cohen’s $\kappa$ is another popular measure but is applicable to only two raters. We calculated Cohen’s $\kappa$ for each human and model pairs. Although we observe lower values with Cohen’s $\kappa$, the difference between human-human and human-model agreements remains similar. We share the detailed results in the supplementary material. We proposed a 0.50 inter-model threshold based on our experiments; this may not always work (e.g., in scenarios where typical agreement values are higher or lower). Another possibility (besides using an inter-model agreement threshold), is to manually label a subset of the data; if human-human and human-model IRAs, are similar in this subset, the model could replace one human.
Few-shot learning requires some human effort initially, to get the “shots”. However, labeling 3–4 samples is much cheaper compared to labeling all the samples. Also, in few-shot learning, we present the model with one randomly chosen example from each category to prevent bias. However, outcomes may depend on the set of few-shot samples used.
Although all of our datasets are relatively recent, some of them might be part of the training data of some of the LLMs we evaluate. As our methodology performs several pre-processing steps, such as function body accumulation (for the semantic similarity task) and diff generation (for the static analysis warnings), it is highly unlikely that the model has memorized the exact results. However, we acknowledge that some bias may exist due to the model’s exposure to the underlying GitHub repositories.
VII Related Work
@@
ML and LLMs in Software Engineering Learning-based approaches and LLMs have been applied to various software development tasks [33], including code completion [34, 35, 36], test generation [37, 38, 39, 40], fuzzing [41, 42], automated repair [43, 44, 45, 46, 47] and coding agents [48, 49, 50, 51], agents for other software engineering tasks [52], and type prediction [53, 54, 55, 56]. We follow this trend, but deviate by not targeting a specific task, but by studying the potential for automating human-subject (manual) evaluations, using LLMs.
@@
LLMs to Complement or Replace Human Annotators Recent and concurrent work explores the idea of complementing or replacing human annotators with LLMs in domains beyond software engineering. Work in natural language processing (NLP) compares LLMs and human participants, typically recruited via crowd-working platforms, and report mixed results. While some studies show that adding LLM labels can improve the aggregated labels of a dataset [57] and sometimes outperform crowd-workers [58], others note that LLMs fail to accurately represent differences between demographic groups [59]. Others suggest to use LLMs to perform sub-tasks in a crowd-sourcing pipeline [60]. Bavaresco et al. [13] propose a benchmark of NLP datasets to compare human and LLM performance. We envision our work to serve as such a benchmark in software engineering. To the best of our knowledge, we are the first to carefully evaluate automated LLM-annotations for software engineering artifacts, and propose actionable guidelines for when and how to use LLMs to create annotations.
@@
Collaborative Human-LLM Annotation Our suggested workflow (Section V) relates to work on human-LLM collaboration for data labeling [61]. Some papers [62, 12] suggest first querying an LLM, and asking humans to refine some of the LLM-provided labels. Instead, we find that inter-model agreement and LLM output probabilities provide an effective way of deciding which labeling tasks can be safely delegated to a model.
@@
LLMs for Data Synthesis and Assessment Beyond using LLMs to partially automate evaluations that would otherwise be performed only by humans, others propose to use LLMs to synthesize additional data to train or fine-tune models for text annotation tasks [63, 64]. There is a recent related survey [65]. A key difference from our work is that training and fine-tuning datasets must be large-scale, but some amount of noise is acceptable, whereas annotations in human-subject studies are typically of smaller scale, but should have high confidence. LLMs have also been proposed as judges of outputs generated by other LLMs, either with a single LLM as the judge [66] or with a set of LLMs that discuss until reaching an agreement [67]. These efforts are primarily about determining human preference, i.e., tasks where different humans may legitimately disagree, whereas many annotation tasks in software engineering have an objectively correct answer that humans can eventually agree upon.
VIII Conclusion
@@
In this paper, we investigate if LLM responses can substitute for human raters in software engineering annotation tasks, and how such substitutions may affect inter-rater agreement. We find that human-model agreements can be fairly consistent with human-agreements, for the majority of the settings we studied. We also find model-model agreement is a good indication of human-model agreement, suggesting that when powerful models trained on giant human corpora agree with each other, they tend to agree also with humans. Finally, we find that an LLM’s confidence (output probability) for a given sample output is a good indication of whether LLM output agrees with majority human rating, for that sample. We caution that the paper considers only discrete (multiple choice) LLM responses, not free-form annotations; furthermore, we have not studied issues of model output bias, nor issues of demographics.
Our scripts and datasets are publicly available: https://zenodo.org/doi/10.5281/zenodo.13146386
Acknowledgments
This work was partially supported by the National Science Foundation under CISE SHF MEDIUM 2107592, the European Research Council (ERC, grant agreements 851895 and 101155832) and by the German Research Foundation within the ConcSys, DeMoCo, and QPTest projects.